OpenStudy (anonymous):

ln (x2-1/x3), x>1 the x2 is x square, the x3 is x times itself 3x

OpenStudy (anonymous):

2ln(x)-[ln(1)-3ln(x)] = 2ln(x) + 3ln(x) = 5ln(x)

OpenStudy (anonymous):

\[\ln 6/\sqrt{x2}+1\]